# Detecting Hidden Triggers: Mapping Non-Markov Reward Functions to Markov

### v2

@@ 4.2 Formulation
Consider that for some $(s,a,s^{\prime})\in(S,A,S)_{o}$ and for some fixed $i,j\in\{1,2,...,K\}$ that when we constrain the sum $\forall r\in r(s,a,s^{\prime})$ to be less than or equal to 1, we limit $\delta_{r}(u_{i},L(s,a,s^{\prime}),u_{j})$ to have at most one outcome, thus resolving our reward conflicts. Importantly, by Equation (10), we encode rewards as conditionals for inferring the structure of the RM. Therefore, if upstream rewards form dependencies with downstream rewards, rewards serve as memory, hence $H=(S\times A\times\mathbb{R})$. Like in Equation (9), Equation (10) would be an equality when spanning $\mathbb{R}$ and $S\times A\times S$ instead of some finite observed space. Regardless, by Equation (10) we satisfy the determinism of $\delta_{r}$ and, thereby, Property 4 of the ARMDP.
Eqs (6-10) form the entirety of the constraints for ARMDP learning. However, there are potentially many ARMDP solutions for some given $T_{o}$. As such, we are motivated to derive simple and interpretable models for representing reward dynamics. Therefore, we orient our objective and learning procedure towards finding the most minimal ARMDP, and thus the most minimal RM, that satisfies $T_{o}$. This is done in two ways. First, for any inference of an ARMDP, we start by assuming $K=|U|=2$, thus dramatically limiting our search space. If $K=2$ is determined to be insufficient to resolve reward conflicts (ILP infeasibility) we increment $K$ by 1 and try again. This is known as a deepening strategy employed in various related works [9, 6]. Second, we limit the number of transition tuples that trigger transitions between RM states (e.g., $\delta_{u}(u_{i},L(s,a,s^{\prime}))=u_{j}$ where $u_{i}\neq u_{j}$). We do so by minimizing the following objective:
@@
We solve Eqs (6-11) using out-of-the-box ILP solvers in Gurobi [14] to infer viable ARMDPs. We note that all notation in this section was constructed assuming $L:S\times A\times S\rightarrow 2^{P}$ though it is straightforward to rewrite it as being sensitive to $L:S\rightarrow 2^{P}$. We provide the full ILP in Appendix A.
We also note a structural assumption that we make in later experimentation. That is, we only use $o_{m,n,i,j}$ where $j\geq i$ (i.e., the upper triangle matrix of each $O_{m,n}$). The implication of this is that we assume the RM to never return to a previous state, once left. While this might seem to limit the expressiveness of any resulting ARMDP we note that our ILP simply unrolls these cycles, under this assumption, and expands $U$ as needed. Because $T_{o}$ represents a finite set of finite trajectories, learned ARMDPs will be equivalently expressive, but larger in some cases. However, for all RMs considered in this work (e.g., the RMs in Figure 1) this structural assumption holds and we still learn the minimal ARMDP while dramatically reducing learning time. We provide more details on this assumption in Appendix B.
@@ 4.3 Extracting RMs from ARMDPs
We provide a brief outline for extracting an RM from an ARMDP. Given that $K=|U|$ is determined by our deepening strategy, we assume $U$ is given. Additionally, by Equation (7) we designate $u_{1}\in U$, to be the initial state. To extract the rules of $\delta_{u}$ and $\delta_{r}$ we simply iterate $\forall m=1,2,...M,\forall n=1,2,...,l_{m},\forall i=1,2,...,K,\forall j=1,2,...,K$. If $O_{m,n,i,j}=1$ then $\delta_{u}(u_{i},L(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}))=u_{j}$ and $\delta_{r}(u_{i},L(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}),u_{j})=r_{n+1}^{m}$.
@@ 4.4 Theoretical Considerations
Our ILP is constructed such that it infers an ARMDP that satisfies reward conflicts in $T_{o}$. However, because the ARMDP is simply just an MDP, we can apply out-of-the-box RL algorithms to solve it. However, it remains to be shown how the ARMDP might be used as a proxy of the NMRDP for the purpose of maximizing reward expectations. Because we assume the form of $R$ in the NMRDP to be an RM, then any observed reward in $T_{o}$ would be consistently reproducible assuming identical histories. Therefore, so long as the likelihood of producing trajectories under $T$ is identical to $\tilde{T}$, reward expectations are preserved. We start with the weighted reward sum for some $\tau_{m}\in T_{o}$:
$R(\tau_{m})=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ (12)
Similarly, we construct a weighted reward sum for some $\tilde{\tau}_{m}$ derived by mapping $\tau_{m}$ onto an ARMDP to be:
$R(\tilde{\tau}_{m})=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$ (13)
We remind readers that every $O_{m,n,i,j}$ reflects a binary variable that, if 1, allows us to conclude $\delta_{u}(u_{i},L(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}))=u_{j}$. By referencing Property 3 of the ARMDP, the interpretation of multiplying $O_{m,n,i,j}$ into $T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ to determine $\tilde{T}$ should be evident. With this we introduce Theorem 1:
@@
$R(\tau_{m})=R(\tilde{\tau}_{m})$ $\forall\tau_{m}\in T_{o}$ where $\tilde{\tau}_{m}$ is derived by mapping $\tau_{m}$ onto an ARMDP.
For space reasons, we provide the proof of Theorem 1 in Appendix C, but state that by Theorem 1, the ARMDP is a suitable proxy for maximizing reward sums $\forall\tau_{m}\in T_{o}$.
@@ 4.5 RL for Non-Markov Rewards
Our ILP formulation for learning ARMDPs is a form of passive learning. That is to say, the ARMDP will be sufficient for $T_{o}$, but may not reflect the ground truth. However, by Theorem 1, we know that we can use ARMDPs as a proxy for maximizing reward expectations for the NMRDP. Therefore, we apply Q-Learning to make our approach an active learner. Our intuition is that we might resolve the reward conflicts that are needed to maximize our expected reward.
@@
6:     $\tau_{m},conflict=NMRDP.sim(ARMDP,\pi_{ARMDP})$
7:     if $conflict$ then
@@
9:        $solved=False$
10:        while $!solved$ do
11:           $ARMDP,solved=solve\_ILP(T_{o},K)$
@@
We present ARMDPQ-Learning, with pseudocode available in Algorithm 1. The agent applies our deepening strategy by initially assuming the problem is an MDP (e.g., $|U|=1$). We denote this ARMDP as ARMDP0 (line 1). $K=|U|=2$, is used for subsequent ARMDP inference. The agent policy, $\pi_{ARMDP}$, is set to $\pi_{0}$, a policy with all Q-values set to 0 (line 2). We use the NMRDP to simulate some $\tau_{m}$ under the perspective of ARMDP, selecting actions according to $\pi_{ARMDP}$ (line 6). If a conflict is found we add $\tau_{m}$ into $T_{o}$ (line 8) and solve for a new ARMDP (line 11). If the problem is infeasible we increment $K$ by 1 (line 13) until there is a solution. As each ARMDP presents a new perspective, we reset $\pi_{ARMDP}$ back to $\pi_{0}$ (line 16). Assuming no conflict is found for $\tau_{m}$, we update $\pi_{ARMDP}$ with $\tau_{m}$ according to Q-Learning updating (line 18). Note that line 18 is shorthand. In reality, tuples are updated in an online fashion while simulating the trajectory. Also, by only adding $\tau_{m}$ known to conflict with our current ARMDP, we keep $T_{o}$ small to make our ILP solving more efficient.
@@ 5.1 Experiment 1 - Officeworld
We validate our approach by applying ARMDPQ-Learning on the four black-box RMs (Tasks (b-e)) from the Officeworld domain visualized in Figure 1. We extend each RM with a terminal state on all * tiles, resulting in a Markov reward of -1. The RM, as well as $P$ and $L$ are hidden from the agent. We refer readers to Appendix D for details on hyper-parameter selection, but note that we use an $\epsilon$-greedy policy and set $\epsilon$ to 0% after 90,000 episodes to exploit learned behavior. Results for Experiment 1 can be seen in Figure 2.
Figure 2: ARMDPQ-Learning evaluated on Tasks (b-e) in the Officeworld domain on 10 random trials. Rewards show +/- a standard deviation.
The agent learns to perform optimally in all tasks, suggesting that a meaningful representation was learned. Task (b), (c), and (e) all arrived at $|U|$ reflecting the ground truth model (2, 2, and 4, respectively). Conversely, Task (d) arrived at $|U|\approx 3.2$. Because the coffee and the mail tiles can be achieved in any order, occasionally the agent would lock on to one solution, and not the other, requiring less RM states. $T_{o}$ remained relatively small for all tasks, with $|T_{o}|\approx$ 4, 8, 39, and 44 for Tasks (b-e), respectively. We also tracked the cumulative solve times for each task to be $\approx$ 0.1, 0.6, 4,200, and 12,000 seconds, respectively. We include an extension to Figure 2 in Appendix D, containing figures for $|U|$, $|T_{o}|$ and cumulative solve times.
In all experiments, we noticed high variability in early performance reflected in high standard deviations between episodes 0-20,000. We attribute this to identifying the wrong pattern in $H$ during early conflict resolution. Interestingly, because these early solutions are so poor, they are easily testable by the Q-Learing agent. This lends some credence to our initial intuition of the Q-Learning agent serving as an important active learner capable of interrogating its own perspectives and updating in the face of new evidence.
@@ 5.2 Experiment 2 - ARMDP Representations
Previously, Christoffersen et al. compared the learning rates of Q-Learning agents over NMRDPs, using DFAs, to sparse, recurrent deep RL models [9]. However, in smaller discrete state spaces, Q-Learning has the advantage of updating state-action values directly given their simpler, tabular representation. In contrast, deep RL models must learn approximate solutions using complex computations over large batches of data. It remains to be shown how beneficial it could be to represent the state space of a DQN with the state space of an ARMDP. This would allow for a more even comparison in learning performance against RDQN architectures that have to learn a Markov representation over the sparse histories they encode.
For each Task (b-e), we extract the ARMDP representations from Experiment 1 and represent $S$ and $U$ as the one-hot encoded feature space of an Abstracted DQN (ADQN). As the ADQN explores each task we update $S$ and $U$ according to its respective ARMDP. In parallel, we train recurrency-based models without access to these abstractions, namely, a Recurrent DQN (RDQN), a Long-Short Term Memory DQN (LSTMDQN), and a Gated Recurrent DQN (GRUDQN). For a baseline, we also train a DQN without any knowledge of $U$.
Figure 3: DQN models evaluated on Tasks (b-e) from the Officeworld domain on 10 random trials.
Rewards are very sparse for the Officeworld tasks and exploration via an $\epsilon$-greedy policy can often be insufficient. As we don’t seek to demonstrate the efficacy of RL strategies under sparse rewards, we occasionally seed each model with an optimal trajectory (every 10th trajectory). Results for Experiment 2 can be seen in Figure 3.
In all tasks, the ADQN model learns at the rate of, or out-learns, all other models. This becomes more evident as the complexity of the task increases (Task (e) being the most complex). We note that for Task (b), the baseline DQN performs just as well as the other models only because the shortest path to the goal includes the coffee tile. Interestingly, for Task (e), the RDQN performs just as poorly as the baseline DQN model while the LSTMDQN and GRUDQN models (while learning slowly) appear to be uncovering the relevant hidden mechanisms of the underlying RM. We hypothesize that the forgetting component of the GRUDQN and LSTDQN is a key differentiating factor in their learning ability. Importantly, we remind readers to still consider the cumulative solve times of ARMDQ-Learning from Experiment 1 as this is a non-trivial time cost.
@@ 5.3 Experiment 3 - Breakfastworld
To evaluate the effectiveness of learning our RMs over $H=(S\times A\times\mathbb{R})^{*}$, rather than $H=(S\times A)^{*}$, we introduce a new domain, Breakfastworld, seen in Figure 4 (a). We include two tasks, Tasks (b) and (c), containing multiple interdependent reward signals. The domain uses $P=\{c,w,e,l\}$ representing propositions for cooking, washing, eating and leaving, respectively. Task (b) contains an RM wherein the agent is meant to cook, eat, and then leave. $(u_{1},c,u_{2})$ returns a reward of -0.1, reflecting a time cost. However, this time cost is experienced in a non-linear fashion as the agent grows bored and spending another turn cooking returns a reward of -0.3, $(u_{2},c,u_{3})$. However, this patience pays off in $(u_{3},e,u_{4})$, where eating a perfectly cooked meal returns a reward of 2. In contrast, $(u_{2},e,u_{5})$ returns 1, reflecting an undercooked meal. Finally, $(u_{4},l,term)$ and $(u_{5},l,term)$ reflect the last reward dependency, returning the cumulative reward for each path in the RM. Task (c) builds on Task (b) by adding a reward dependency for washing. The penalty for washing after cooking once, $(u_{6},w,u_{7})$, is less than after cooking twice, $(u_{4},w,u_{5})$.
@@
Both tasks reflect RMs more complex in terms of $|U|$ than Task (e) in Experiment 1. However, we will demonstrate that RM learning benefits in cases where rewards are interdependent, as observed reward conflicts result in natural splits in the ARMDP learning by Equation (10). Conversely, DFA learning suffers under such conditions as each reward conflict must be represented by its own DFA [13]. Moreover, downstream DFAs are learned independently from upstream DFAs as they only learn according to $H=(S\times A)^{*}$.
We demonstrate how advantageous it can be that our ILP naturally learns according to $H=(S\times A\times\mathbb{R})^{*}$ by applying ARMDPQ-Learning to Tasks (b) and (c). We then compare this to applying ARMDPQ-Learning on RMs equivalent in structure to Task (b) and (c), but that only emit the cumulative rewards on the terminal state. All other rewards are made 0. We refer to Tasks (b) and (c), in Figure 4, as Task (b) Full and Task (c) Full, respectively. Conversely, the RMs that only emit the cumulative reward are referred to as Task (b) Cumulative and Task (c) Cumulative, respectively. The purpose of these cumulative tasks is to form a problem that produces reward conflicts on tile [0,0] that can not be contextualized by the prior reward conflicts. This would be similar to learning two DFAs to parse the two different possible rewards on tile [0,0] only using $H=(S\times A)^{*}$. We show results for Experiment 3 in Table 1.
In all solvable cases of ARMDP-Learning, the agent achieved optimal performance and identified the correct $|U|$ for each task. We include an extension figure to Table 1 in Appendix F containing such results. Using Task (b) Full our cumulative solve time ($\approx$ 4.4 seconds) was dramatically faster than when using Task (b) Cumulative ($\approx$ 1,102.9 seconds). This can be attributed to Task (b) Full requiring roughly a fourth of the number of trajectories to solve than Task (b) Cumulative, likely because of the leveraged context of upstream rewards. More impressive were the results for Task (c) Full which resulted in the cumulative solve time of $\approx$ 153.8 seconds. In no instance of attempting to learn in Task (c) Cumulative did we arrive at a stable solution in 100,000 seconds ($>$27 hours).

### v3

@@ 4.2 Formulation
Consider that for any $(s,a,s^{\prime})\in(S,A,S)_{o}$ and for any $i,j\in\{1,2,...,K\}$ that when we constrain the sum $\forall r\in r(s,a,s^{\prime})$, in Equation (10), to be less than or equal to 1, we limit $\delta_{r}(u_{i},L(s,a,s^{\prime}),u_{j})$ to have at most one reward outcome, thus resolving our reward conflicts in $T_{o}$. Importantly, $I_{s,a,r,s^{\prime},i,j}=1$ contextualizes its corresponding hidden trigger, $I_{s,a,s^{\prime},i,j}$, with reward. When rewards are interdependent (e.g., some upstream reward conflict determines a downstream reward conflict), rewards serve as memory, hence, we capture relationships in $H=(S\times A\times\mathbb{R})^{*}$. By Equation (10), we satisfy the determinism of $\delta_{r}$ and Property 4 of the ARMDP.
Equations (6-10) form the entirety of the constraints for ARMDP learning. However, there are potentially many ARMDP solutions for any given $T_{o}$. As such, we are motivated to derive simple and interpretable models for representing reward dynamics. Therefore, we orient our objective and learning procedure toward finding the minimal ARMDP, and thus the minimal RM, that satisfies $T_{o}$. This is done in two ways. First, for any inference of an ARMDP, we start by assuming $K=|U|=2$, thus dramatically limiting our search space. If $K=2$ is determined to be insufficient to resolve reward conflicts (ILP infeasibility) we increment $K$ by 1 and try again. This is known as a deepening strategy employed in various related works [9, 6]. Second, we limit the number of hidden triggers that dictate transitions between different RM states (i.e., $\delta_{u}(u_{i},L(s,a,s^{\prime}))=u_{j}$ where $u_{i}\neq u_{j}$). We do so by minimizing the following objective:
@@
We solve Equations (6-11) using out-of-the-box ILP solvers in Gurobi [13] to infer viable ARMDPs. We note that all notation in this section was constructed assuming $L:S\times A\times S\rightarrow 2^{P}$ though it is straightforward to rewrite it as being sensitive to $L:S\rightarrow 2^{P}$. We provide the full ILP in Appendix A.
We also note a structural assumption that we make in later experimentation. That is, we only use $O_{m,n,i,j}$ where $j\geq i$ (i.e., the upper triangle matrix of each $O_{m,n}$). The implication of this is that we assume the RM to never return to a previous state, once left. While this might seem to limit the expressiveness of any resulting ARMDP, we note that our ILP simply unrolls these cycles expanding $U$ as needed. Because $T_{o}$ represents a finite set of finite length trajectories, learned ARMDPs will be equivalently expressive, but larger in some cases. For all RMs considered in this work, this structural assumption holds resulting in the minimal ARMDP being learned while dramatically reducing learning time. We provide more details on this assumption in Appendix B.
@@ 4.3 Extracting RMs from ARMDPs
We provide a brief outline for extracting an RM from an ARMDP. Given that $K=|U|$ is determined by our deepening strategy, we assume $U$ is given. Additionally, by Equation (7) we designate $u_{1}\in U$ to be the initial state. To extract the rules of $\delta_{u}$ and $\delta_{r}$ we simply iterate $\forall m=1,2,...M,\forall n=1,2,...,l_{m},\forall i=1,2,...,K,\forall j=1,2,...,K$. If $O_{m,n,i,j}=1$ then $\delta_{u}(u_{i},L(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}))=u_{j}$ and $\delta_{r}(u_{i},L(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}),u_{j})=r_{n+1}^{m}$.
@@ 4.4 Theoretical Considerations
Our ILP is constructed such that it infers an ARMDP that resolves reward conflicts in $T_{o}$. Because the ARMDP is simply just an MDP, we can apply out-of-the-box RL algorithms to solve it. However, it remains to be shown how the ARMDP might be a suitable proxy for the NMRDP for the purpose of maximizing reward expectations. We start with the weighted Reward Sum (RS) for some $\tau_{m}\in T_{o}$:
$RS(\tau_{m})=\sum_{n=1}^{l_{m}}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ (12)
Given that we assume the form of $R$ to be an RM and that $\delta_{u}$ and $\delta_{r}$, in an RM, are deterministic functions, $RS(\tau_{m})$ reliably captures the reward observations essential for constructing a reward expectation, assuming the trajectory sequence is preserved. Similarly, we construct a weighted Abstracted Reward Sum (ARS) for some $\tau_{m}\in T_{o}$ under the representation of the ARMDP to be:
$ARS(\tau_{m})=\sum_{n=1}^{l_{m}}\sum_{i=1}^{K}\sum_{j=1}^{K}r_{n+1}^{m}\cdot T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})\cdot O_{m,n,i,j}$ (13)
We remind readers that every $O_{m,n,i,j}$ reflects a binary variable that, if 1, indicates $\delta_{u}(u_{i},L(s_{n}^{m},a_{n}^{m},s_{n+1}^{m}))=u_{j}$. By referencing Property 3 of the ARMDP, multiplying $O_{m,n,i,j}$ into $T(s_{n}^{m},a_{n}^{m},s_{n+1}^{m})$ to determine $\tilde{T}$ should be evident. We introduce Theorem 1:
@@
$RS(\tau_{m})=ARS(\tau_{m})$ $\forall\tau_{m}\in T_{o}$.
For space reasons, we provide the proof of Theorem 1 in Appendix C, but briefly state that by ensuring $RS(\tau_{m})=ARS(\tau_{m})$ $\forall\tau_{m}\in T_{o}$, the ARMDP is a suitable proxy for maximizing reward expectations for the NMRDP, assuming $T_{o}$ is representative. Furthermore, even if $T_{o}$ is not representative, we posit that any ARMDP derived from $T_{o}$ serves as a useful hypothesis for RL agents to interrogate in an active learning paradigm. We explore that intuition in the next section.
@@ 4.5 RL for Non-Markov Rewards
Our ILP formulation for learning ARMDPs is a form of passive learning. That is to say, while the ARMDP will sufficiently resolve reward conflicts found in $T_{o}$, $T_{o}$ may not be representative of the true reward dynamics. Therefore, we are interested in extending our approach into an active learning framework by applying Q-Learning over inferred ARMDPs. Our intuition is that while any instance of $T_{o}$, or any ARMDP inferred from it, may not be truly representative, Q-Learning agents might gradually expand $T_{o}$ in their pursuit to maximize reward expectations. In this way, each inferred ARMDP acts as a hypothesis that can be interrogated for consistency.
@@
[…]
@@
[…]
@@
[…]
@@ 5.1 Experiment 1 - Officeworld
We validated our approach by applying ARMDPQ-Learning on the four black-box RMs (Tasks (b-e)) from the Officeworld domain visualized in Figure 1. We hid the RM as well as $P$ and $L$ from the agent. We refer readers to Appendix D for details on hyper-parameter selection, but note that we used an $\epsilon$-greedy policy and set $\epsilon$ to 0% after 90,000 episodes to exploit learned behavior. Results for Experiment 1 can be seen in Figure 2.
Figure 2: Average rewards with standard deviations, evaluated over 10 trials using ARMDPQ-Learning on Tasks (b-e) in the Officeworld domain.
The agent learned to perform optimally in all tasks, suggesting that a meaningful representation was learned. Task (b), (c), and (e) all arrived at $|U|$ reflecting the ground truth model (2, 2, and 4, respectively). Conversely, Task (d) arrived at $|U|\approx 3.2$. Because the coffee and the mail tiles can be achieved in any order, occasionally the agent would lock on to one solution, and not the other, requiring less RM states. $T_{o}$ remained relatively small for all tasks, with $|T_{o}|\approx$ 4, 8, 39, and 44 for Tasks (b-e), respectively. We also tracked the cumulative solve times for all ILP solves under each task to be $\approx$ 0.1, 0.6, 4,200, and 12,000 seconds, respectively. We include an extension to Figure 2 in Appendix D, containing figures for $|U|$, $|T_{o}|$ and cumulative solve times.
In all experiments, we noticed high variability in early performance reflected in high standard deviations between episodes 0-20,000. We attribute this to identifying the wrong pattern in $H$ during early conflict resolution. Interestingly, because these early solutions were so poor, they were easily testable by the Q-Learning agent. This lends some credence to our initial intuition of the Q-Learning agent serving as an important active learner capable of interrogating its own perspectives and updating in the face of new evidence.
@@ 5.2 Experiment 2 - ARMDP Representations
Previously, Christoffersen et al. compared the learning rates of Q-Learning agents over NMRDPs, using DFAs, to sparse, recurrent, deep RL models [9]. However, in smaller discrete state spaces, Q-Learning has the advantage of updating state-action values directly given their simpler, tabular representation. In contrast, deep RL models must learn approximate solutions using complex computations over large batches of data. It remained to be shown how beneficial it could be to represent the state space of a DQN with $\tilde{S}$ instead of $S$. This would allow for a more even comparison of training profiles against RDQN architectures that have to learn a Markov representation over histories of $S$.
For each Task (b-e), in Experiment 1, we extracted $\tilde{S}$, from the finalized ARMDP. We used $\tilde{S}$ as the feature space for an Abstracted DQN (ADQN). As the ADQN explored each task we updated $\tilde{S}$ according to its respective ARMDP. In parallel, we trained recurrency-based models using $S$, namely, an RDQN, a Long-Short Term Memory DQN (LSTMDQN), and a Gated Recurrent DQN (GRUDQN). For a baseline, we also trained a DQN using $S$.
Figure 3: Average rewards with standard deviations, evaluated over 10 trials using DQN models on Tasks (b-e) in the Officeworld domain.
Rewards were very sparse for the Officeworld tasks and exploration via an $\epsilon$-greedy policy was often insufficient. As we didn’t seek to demonstrate the efficacy of RL strategies under sparse rewards, we seeded each model with an optimal trajectory every 10th trajectory. Results for Experiment 2 can be seen in Figure 3.
In all tasks, the ADQN model learned at the rate of, or out-learned, all other models. This became more apparent as the complexity of the task increased (Task (e) being the most complex). We note that for Task (b), the baseline DQN performed just as well as the other models only because the shortest path to the goal included the coffee tile. Interestingly, for Task (e), the RDQN performed just as poorly as the baseline DQN model while the LSTMDQN and GRUDQN models (while learning slowly) appeared to be uncovering the relevant hidden mechanisms of the underlying RM. We hypothesize that the forgetting component of the GRUDQN and LSTDQN is a key differentiating factor in their learning ability. Still, we remind readers that uncovering $U$, as discussed in Experiment 1, involves a non-trivial cost before ADQN training can commence.
@@ 5.3 Experiment 3 - Breakfastworld
To evaluate the effectiveness of learning our RMs over $H=(S\times A\times\mathbb{R})^{*}$, rather than $H=(S\times A)^{*}$, we introduce a new domain, Breakfastworld, seen in Figure 4 (a). We include two tasks, Tasks (b) and (c), containing multiple interdependent reward signals. The domain uses $P=\{c,w,e,l\}$ representing propositions for cooking, washing, eating and leaving, respectively. Task (b) contains an RM wherein the agent is meant to cook, eat, and then leave. $(u_{1},c,u_{2})$ returns a reward of -0.1, reflecting a time cost. This time cost is experienced in a non-linear fashion as the agent grows bored and spending another turn cooking returns a reward of -0.3, $(u_{2},c,u_{3})$. However, this patience pays off in $(u_{3},e,u_{4})$, where eating a well cooked meal returns a reward of 2. In contrast, $(u_{2},e,u_{5})$ returns 1, reflecting an undercooked meal. Finally, $(u_{4},l,term)$ and $(u_{5},l,term)$ reflect the last reward dependency, returning the cumulative reward for each path in the RM, encouraging episode completion. Task (c) builds on Task (b) by adding a reward dependency for washing. The penalty for washing after cooking once, $(u_{6},w,u_{7})$, is less than after cooking twice, $(u_{4},w,u_{5})$.
@@
Both tasks reflect RMs more complex in terms of $|U|$ than Task (e) in Experiment 1. However, we demonstrate that RM learning benefits in cases where rewards are interdependent as upstream reward conflict resolution results in natural splits for downstream reward conflict resolution. Conversely, DFA learning suffers under such conditions as each reward conflict must be represented by its own DFA which is learned independently from any other DFA [12], hence they only learn according to $H=(S\times A)^{*}$.
We demonstrate how advantageous it can be that our ILP naturally learns according to $H=(S\times A\times\mathbb{R})^{*}$ by applying ARMDPQ-Learning to Tasks (b) and (c). We then compare this to applying ARMDPQ-Learning on RMs equivalent in structure to Task (b) and (c), but that only emit cumulative rewards on terminal transitions. All other rewards are made 0. We refer to Tasks (b) and (c), in Figure 4, as Task (b) Full and Task (c) Full, respectively. Conversely, the RMs that only emit the cumulative rewards are referred to as Task (b) Cumulative and Task (c) Cumulative, respectively. The purpose of these cumulative tasks is to form a problem that produces reward conflicts only on the $l$ tile that can not be contextualized by upstream reward conflicts. This would be similar to learning DFAs to parse the different possible rewards on the $l$ tile only using $H=(S\times A)^{*}$. We show results for Experiment 3 in Table 1.
In all solvable cases of ARMDPQ-Learning, the agent achieved optimal performance and identified the correct $|U|$, for each task. We include an extension figure to Table 1 in Appendix F containing such results. Using Task (b) Full our cumulative solve time for all ILPs was dramatically smaller ($\approx$ 4.4 seconds) than when using Task (b) Cumulative ($\approx$ 1,102.9 seconds). This can be attributed to Task (b) Full requiring only a fourth of the number of trajectories to solve than Task (b) Cumulative, likely due to the leveraged context of upstream rewards. More impressive were the results for Task (c) Full which resulted in the cumulative solve time of $\approx$ 153.8 seconds. In no instance of attempting to learn in Task (c) Cumulative did we arrive at a stable solution in 100,000 seconds ($>$27 hours).